OpenStudy (anonymous):

what are the real or imaginary solutions of the polynomial equation x4 – 41x2 = –400

hero (hero):

let y = x^2 y^2 - 41y = -400 y^2 - 41y + 420.25 = -400 + 420.25 (y - 20.5)^2 = 20.25 y - 20.5 = ±sqrt(20.25) y =±sqrt(20.25) + 20.5 substitute x^2 back in: x^2 = ±sqrt(20.25) + 20.5 x = ±sqrt(±sqrt(20.25) + 20.5) Believe it or not, that leads to four answers

OpenStudy (anonymous):

can you help me with the rest of my math ? im extremely far behind.

hero (hero):

I do not have time to do that, sorry.

OpenStudy (anonymous):

okayy well ... explain this one too me..

hero (hero):

Basically, I completed the square to solve it.